OpenStudy (anonymous):

26xy - 39 divided by 13y

OpenStudy (klimenkov):

\(2x-\frac3y\)

OpenStudy (anonymous):

Is this the expression?\[(26x y-39 )/ 13y \]

OpenStudy (anonymous):

\[\frac{ 26xy }{ 13y } - \frac{ 36 }{ 13y }\] \[\frac{ 2x }{ 1 } - \frac{ 3 }{ y }\]

OpenStudy (anonymous):

\[(26x y-39 )/ 13y=2 x y^2-3 y \]\[(26x y-39 )/(13y)=2 x-\frac{3}{y} \]